OpenStudy (anonymous):

Which of the following integers must be a factor of the sum of any 3 consecutive positive integers? 2 3 4 5 6

OpenStudy (anonymous):

say the number is n+(n+1)+(n+2)=3n+3=3(n+1) see the answer?

OpenStudy (anonymous):

yes thank you